# Deep Learning for Clouds and Cloud Shadow Segmentation in Methane Satellite and Airborne Imaging Spectroscopy

### v2

@@ IV-B5 Combined Models
This model fuses U-Net and SCAN predictions using a multilayer perceptron (MLP). The architecture employs two pre-trained base models (U-Net and SCAN) with frozen weights to maintain their individual predictive capabilities. Given the input hyperspectral image $X\in\mathbb{R}^{B\times H\times W\times C}$, the predictions from the base models are:
$P_{U-Net}(y|X)\in\mathbb{R}^{B\times H\times W\times K}$
$P_{SCAN}(y|X)\in\mathbb{R}^{B\times H\times W\times K}$
where $K$ is the number of classes. These predictions are concatenated channel-wise to form the input to the merging MLP:
$P_{combined}=\text{Concat}[P_{U-Net},P_{SCAN}]\in\mathbb{R}^{B\times H\times W\times 2K}$
The MLP fusion network with $M$ hidden layers processes each pixel independently:
$f_{fusion}(p)$ $=W_{M+1}(\phi(BN_{M}(W_{M}($
$\quad\phi(BN_{1}(W_{1}P_{combined}))))))$ (1)
where $W_{i}$ are learnable weight matrices, $BN_{i}$ denotes batch normalization, and $\phi$ is the ReLU activation function. Dropout with rate $\delta=0.2$ is applied after each hidden layer for regularization. The final output provides the fused class probabilities as follows:
$P(y|X)=\text{Softmax}(f_{fusion}(P_{combined}))\in\mathbb{R}^{K}$
@@
This approach employs convolutional layers to merge predictions, preserving spatial context during fusion. Using the same pre-trained U-Net and SCAN models with frozen weights, the concatenated predictions $P_{combined}\in\mathbb{R}^{B\times H\times W\times 2K}$ are processed through a series of convolutional layers:
$F_{1}=\phi(BN_{1}(\text{Conv}_{1}(P_{combined})))\in\mathbb{R}^{B\times H\times W\times C_{1}}$
$F_{i}$ $=\phi(BN_{i}(\text{Conv}_{i}(F_{i-1})))\in\mathbb{R}^{B\times H\times W\times C_{i}},$
@@
where $\text{Conv}_{i}$ represents a 2D convolutional layer with kernel size $3\times 3$ and padding size 1 (except for the final layer which uses $1\times 1$ convolution), preserving original spatial dimensions. $BN_{i}$ denotes batch normalization, $\phi$ is the ReLU activation function, and $C_{i}$ are the channel dimensions of the intermediate feature maps (e.g., $C_{1}=64$, $C_{2}=32$, $C_{3}=16$ in our work). Dropout with rate $\delta=0.2$ is applied after each intermediate convolutional layer. The final class probabilities are obtained through:
$P(y|X)=\text{Softmax}(F_{N})$
Unlike the MLP approach, the CNN merger maintains the spatial structure of predictions throughout the fusion process. Each convolutional layer captures local relationships between the predictions of both models, allowing the network to learn spatial patterns in prediction agreement or disagreement. This spatial-aware fusion can be particularly beneficial for complex scenes where the performance of individual models varies spatially.
IV-C Training
@@
$\mathcal{L}=-\sum_{n=1}^{N}\sum_{c=1}^{C}w_{c}y_{n,c}\log(\hat{y}_{n,c})$
where $C$ represents the number of classes, $w_{c}$ denotes the weight assigned to class $c$, $y_{n,c}$ is the ground truth binary indicator (1 if sample $n$ belongs to class $c$, and 0 otherwise), and $\hat{y}_{n,c}$ represents the model’s predicted probability that sample $n$ belongs to class $c$. The class weights $w_{c}$ are computed as the inverse of the class frequencies in the training set, helping to address the inherent imbalance between cloud, shadow, and clear sky spatial soundings.
For optimization, we employ the Adam optimizer with a learning rate determined through cross-validation experiments detailed in Appendix Deep Learning for Clouds and Cloud Shadow Segmentation in Methane Satellite and Airborne Imaging Spectroscopy. To enhance model generalization, we implement data augmentation techniques including random horizontal and vertical flips, and rotations at multiples of 90 degrees. These augmentations help the models learn invariance to common geometric transformations while preserving the physical meaning of the spectral signatures. Training proceeds over a maximum of 100 epochs with randomly shuffled batches of size 32. We employ an early stopping strategy to prevent overfitting, where training is terminated if the validation loss does not improve for 20 consecutive epochs (patience period), and we save the best-performing model checkpoint based on validation performance for final evaluation.
IV-D Post-processing and Evaluation
IV-D1 Post-processing for MethaneSAT
@@ IV-D2 Evaluation
  • Precision: Represents the model’s exactness in identifying clouds and shadows, calculated as the ratio of correctly predicted cloud/shadow spatial soundings to all spatial soundings predicted as cloud/shadow by the model.. High precision indicates a low false positive rate, meaning when the model predicts a pixel as cloud or shadow, it is likely to be correct. Precision = $\frac{\text{True Positives}}{\text{True Positives + False Positives}}$
  • Recall: Measures the completeness of the model in detecting actual clouds and shadows, calculated as the ratio of correctly predicted cloud/shadow pixels to all ground truth cloud/shadow pixels in the dataset. High recall indicates the model’s ability to identify most of the actual clouds and shadows. Recall = $\frac{\text{True Positives}}{\text{True Positives + False Negatives}}$
@@
We use macro-averaging, computing the precision, recall, and F1-score (i.e. we compute the metrics for each class independently and average them to report final results). The reported values include the test set mean performance and standard deviation in 3 cross-validated folds, ensuring a robust evaluation of model performance. See Appendix Deep Learning for Clouds and Cloud Shadow Segmentation in Methane Satellite and Airborne Imaging Spectroscopy for details of the hyperparameter selection.
@@ V-A MethaneAIR
Detailed hyperparameter experiments can be found in Appendix Deep Learning for Clouds and Cloud Shadow Segmentation in Methane Satellite and Airborne Imaging Spectroscopy, and comprehensive predictions for all models can be found in Appendix IV.
@@ VI Discussion
The baseline ILR model (accuracy: 73.81±4.05 for MethaneAIR) captures basic spectral features but generates highly fragmented predictions with significant noise. This purely spectral approach, while computationally efficient, fails to leverage the spatial relationships crucial for distinguishing between spectrally similar features like shadows and dark surfaces. The MLP model (accuracy: 82.49±2.24%) shows improved performance but still suffers from noise and fragmentation due to its pixel-wise classification approach, as clearly visible in Figures 7 and 10.
The U-Net architecture (accuracy: 88.26±0.45% for MethaneAIR, 78.73±3.23% for MethaneSAT) represents a substantial advancement, leveraging its encoder-decoder structure with skip connections to produce spatially coherent predictions. While U-Net exhibits a tendency toward smoothed boundaries around cloud formations and their shadows, this characteristic can be advantageous depending on the specific application requirements. On the other hand, our proposed SCAN model (accuracy: 86.51±2.90% for MethaneAIR, 80.33±3.43% for MethaneSAT) demonstrates superior ability in detecting accurate boundaries through its spectral attention mechanism, though it still exhibits some noise in spectrally complex regions.
Notably, the SCAN model outperforms U-Net on the MethaneSAT data despite showing slightly lower performance on MethaneAIR. This reversal suggests that spectral attention mechanisms may be particularly valuable for the specific spectral characteristics of MethaneSAT observations, highlighting the importance of architecture selection based on the specific instrument’s spectral resolution and characteristics. This finding has significant implications for the development of specialized algorithms for different hyperspectral platforms.
@@
These findings have direct implications for the MethaneSAT mission, which requires accurate identification of atmospheric artifacts to ensure reliable methane retrievals. The superiority of the Combined CNN approach in handling complex scenes with varying surface reflectance is particularly relevant for global methane monitoring, where diverse terrain and atmospheric conditions can significantly impact detection accuracy. By effectively eliminating clouds and shadows from analysis, this approach can substantially improve the reliability of methane concentration mapping, supporting global efforts to identify and quantify methane emissions.
@@ VII Conclusions
From a computational perspective, our analysis demonstrates that the Combined CNN model offers an effective balance between performance and efficiency. While requiring moderately increased computational resources compared to individual models (4.1 ms inference time versus 1.7-2.1 ms per 1,000 km²), its significantly improved segmentation accuracy justifies this overhead for operational deployment in satellite missions where accurate atmospheric artifact detection is critical for reliable methane retrievals and subsequent emission quantification.
The patch-based evaluation strategy demonstrated for MethaneSAT’s variable-sized inputs enhances the practical applicability of these methods in operational contexts. By providing more accurate masks for clouds and shadows, these approaches can significantly improve the reliability of methane concentration mapping, contributing to global efforts to monitor and mitigate methane emissions, a critical component in addressing climate change.
@@ VIII Acknowledgments
Funding for MethaneSAT and MethaneAIR activities was provided in part by Anonymous, Arnold Ventures, The Audacious Project, Ballmer Group, Bezos Earth Fund, The Children’s Investment Fund Foundation, Heising-Simons Family Fund, King Philanthropies, Robertson Foundation, Skyline Foundation and Valhalla Foundation. For a more complete list of funders, please visit www.methanesat.org. We thank the AstroAI and EarthAI institutes at the Center for Astrophyiscs — Harvard & Smithsonian for useful discussions and guidance. CG was supported by AstroAI at the Center for Astrophysics — Harvard and Smithsonian.
Appendix A: Hyperparameter Selection
For hyperparameter optimization, we employed a systematic approach using a 3-fold cross-validation strategy on a randomly selected 10% subsample of the training data.
For each model, we evaluated performance across learning rates $\{1\times 10^{-4},5\times 10^{-4},1\times 10^{-3},5\times 10^{-3},1\times 10^{-2}\}$, using 3-fold cross-validation on the training set. Table IV shows the chosen learning rates for each model architecture for MethaneAIR and MethaneSAT.
For all models, we employed the Adam optimizer with default $\beta_{1}=0.9$ and $\beta_{2}=0.999$ parameters. We applied class weighting to address class imbalance, with weights computed as the inverse of the class frequencies in the training set. No specific normalization was applied beyond the preprocessing steps described in Section IV-A.
The remaining hyperparameters were set according to the architectural specifications detailed in the methodology section. For the MLP, we used hidden layer dimensions of [20, 20]. The U-Net architecture maintained the channel dimensions specified in Section 3.2.3, while the SCAN model used a reduction ratio of 16 for the attention mechanism. The Combined MLP employed hidden dimensions of [256, 128] with a dropout rate of 0.2, and the Combined CNN used channel dimensions of [64, 32, 16] with the same dropout rate.
These hyperparameter configurations were consistently applied across all experimental evaluations to ensure fair comparison between the different architectures.
Model MethaneAIR MethaneSAT
ILR $\times 10^{-2} 1×1$$\times 10^{-2}$
MLP $\times 50^{-3} 1×1$$\times 10^{-2}$
SCAN $\times 10^{-3} 1×1$$\times 10^{-3}$
U-Net $1\times 10^{-3}$ $5\times 10^{-3}$
Combined MLP $1\times 10^{-2}$ $5\times 10^{-4}$
Combined CNN $1\times 10^{-2}$ $5\times 10^{-4}$
TABLE IV: Best learning rates for each data source and model architecture.
Appendix B: Detailed Model Predictions
This appendix provides a comprehensive visual comparison of predictions all evaluated models across both the MethaneAIR and MethaneSAT datasets. These supplementary results validate and expand upon the analyses presented in Section V.
-A MethaneAIR Predictions
Figures 11 through 14 present detailed prediction results for all evaluated models on the MethaneAIR dataset. These visualizations offer insight into the specific capabilities and limitations of each approach when applied to identical input data.
The ILR predictions (Figure 11) exhibit significant fragmentation and noise, particularly evident in regions with complex terrain features. This visual noise corresponds to the high misclassification rates shown in the confusion matrix presented in the main results section, confirming the limitations of purely spectral approaches when spatial context is not considered.
Figure 11: Radiance at 1592nm, ground truth labels, and predictions of the ILR model for MethaneAIR data. The model exhibits significant noise and fragmentation, particularly in regions with complex terrain features.
Figure 12: Radiance at 1592nm, ground truth labels, and predictions of the MLP model for MethaneAIR data. While showing improved coherence compared to ILR, the model still produces speckled predictions.
Figure 13: Radiance at 1592nm, ground truth labels, and predictions of the SCAN model for MethaneAIR data. The model shows enhanced boundary detection capabilities compared to U-Net, but with some residual noise in spectrally complex regions.
Figure 14: Radiance at 1592nm, ground truth labels, and predictions of the Combined MLP model for MethaneAIR data. The fusion approach demonstrates improved performance over individual models.
The MLP model results (Figure 12) show improved coherence compared to ILR but still produce speckled predictions characteristic of pixel-wise classification approaches. While the model better captures the general distribution of clouds and shadows, it struggles with consistent classification, especially in areas with varying surface reflectance.
U-Net predictions (Figure 5) demonstrate a substantial improvement in spatial coherence, with smoother, more contiguous regions for each class. However, as noted in the main results, the model’s convolutional architecture tends to produce over-smoothed boundaries that do not precisely capture the intricate edges of cloud formations and their shadows.
The SCAN model results (Figure 13) illustrate its enhanced boundary detection capabilities compared to U-Net, with more precise delineation of cloud and shadow edges. This improved boundary precision comes with some residual noise in spectrally complex regions, representing the trade-off between spatial coherence and spectral fidelity.
The Combined MLP approach (Figure 14) demonstrates the advantages of ensemble learning, with predictions that balance the strengths of individual models. By adaptively weighting the contributions of both U-Net and SCAN, this fusion approach reduces both the over-smoothing tendency of U-Net and the noise artifacts present in SCAN predictions.
Finally, the Combined CNN results (Figure 4) achieve the most balanced and accurate predictions across all scenes. The model successfully integrates the precise boundary detection of SCAN with the spatial coherence of U-Net, validating the quantitative superiority demonstrated in the confusion matrices presented in the main results section.
-B MethaneSAT Results
Figures 15 through 19 present detailed prediction results for all evaluated models on the MethaneSAT dataset. These visualizations complement the confusion matrices presented in Section V and demonstrate the performance of each model across diverse terrain and atmospheric conditions.
U-Net results (Figure 17) reveal significantly enhanced spatial coherence with the elimination of most noise artifacts. However, as noted in the main results, the model tends to undershoot the extent of cloud formations in certain scenes and produces overly simplified shadow boundaries, limiting its overall accuracy.
The SCAN predictions (Figure 18) demonstrate notably improved performance compared to the U-Net model for MethaneSAT data, particularly in capturing the boundaries of cloud and shadow regions. This visual observation corresponds to the quantitative results presented in the main section, where SCAN outperformed U-Net on MethaneSAT data despite showing slightly lower performance on MethaneAIR.
The Combined MLP approach (Figure 19) shows enhanced performance through fusion, particularly in challenging scenes with varying surface reflectance. The model effectively reduces misclassification between spectrally similar classes while preserving the detailed structures of cloud formations.
Finally, the Combined CNN results (Figure 8) achieve the most accurate and balanced predictions across all MethaneSAT scenes. The model’s ability to preserve spatial context during fusion enables it to effectively handle complex patterns of cloud and shadow formations while minimizing noise and misclassification, validating its superiority as demonstrated in the confusion matrices presented in the main results section.
Figure 15: Data, ground truth labels, and predictions of the ILR model for MethaneSAT data. The predictions exhibit significant artifacts and striping patterns characteristic of sensor-specific noise.
Figure 16: Data, ground truth labels, and predictions of the MLP model for MethaneSAT data. The model shows improved performance over ILR but retains horizontal striping artifacts.
Figure 17: Data, ground truth labels, and predictions of the U-Net model for MethaneSAT data. The model produces spatially coherent predictions but tends to undershoot the extent of cloud formations in certain scenes.
Figure 18: Data, ground truth labels, and predictions of the SCAN model for MethaneSAT data. The model demonstrates improved performance compared to U-Net, particularly in capturing the boundaries of cloud and shadow regions.
Figure 19: Data, ground truth labels, and predictions of the Combined MLP model for MethaneSAT data. The model shows enhanced performance through fusion, particularly in challenging scenes with varying surface reflectance.

### v3

@@ IV-C5 Combined Models
This model fuses model predictions using a multilayer perceptron (MLP). The architecture employs two pre-trained base models (U-Net and SCAN) with frozen weights to maintain their individual predictive capabilities. Given the input hyperspectral image $\mathbf{X}$, we first obtain predictions from frozen pre-trained models:
$\mathbf{P}_{\text{UNet}}$ $=f_{\text{UNet}}(\mathbf{X};\mathbf{W}_{\text{UNet}}^{*})\in\mathbb{R}^{H\times W\times K},$ (11)
$\mathbf{P}_{\text{SCAN}}$ $=f_{\text{SCAN}}(\mathbf{X};\mathbf{W}_{\text{SCAN}}^{*})\in\mathbb{R}^{H\times W\times K},$ (12)
where $\mathbf{W}^{*}$ denotes fixed weights trained independently as described in Sections IV-C3 and IV-C4. Freezing base model weights prevents degradation during fusion training.
We concatenate predictions channel-wise:
$\mathbf{P}_{\text{combined}}=[\mathbf{P}_{\text{UNet}}\,||\,\mathbf{P}_{\text{SCAN}}]\in\mathbb{R}^{H\times W\times 2K}$ (13)
The MLP fusion network with $M$ hidden layers processes each spatial sounding independently:
$f_{fusion}(\mathbf{X})=\mathbf{W}_{M}(\phi(\mathbf{W}_{M-1}(...\phi(\mathbf{W}_{1}P_{combined})))),$
where $\mathbf{W}_{i}$ are learnable weight matrices, and $\phi$ is the ReLU activation function. Dropout with rate $\delta=0.2$ is applied after each hidden layer for regularization. The final output provides the fused class probabilities as follows:
$P(\mathbf{Y}|\mathbf{X})=\text{Softmax}(f_{fusion}(P_{combined}))$
@@
While MLP fusion treats each spatial sounding independently, CNN fusion preserves spatial relationships during combination, enabling the fusion network to leverage spatial context when resolving disagreements between base models. Using the same pre-trained U-Net and SCAN models with frozen weights, the concatenated predictions $P_{combined}\in\mathbb{R}^{B\times H\times W\times 2K}$ are processed through a series of convolutional layers:
$F_{1}=\phi(\text{Conv}_{1}(P_{combined}))\in\mathbb{R}^{B\times H\times W\times C_{1}}$
$F_{i}$ $=\phi(\text{Conv}_{i}(F_{i-1}))\in\mathbb{R}^{B\times H\times W\times C_{i}},$
@@
where $\text{Conv}_{i}$ represents a 2D convolutional layer with kernel size $3\times 3$ and padding size 1 (except for the final layer which uses $1\times 1$ convolution), preserving original spatial dimensions. $\phi$ is the ReLU activation function, and $C_{i}$ are the channel dimensions of the intermediate feature maps (e.g., $C_{1}=64$, $C_{2}=32$, $C_{3}=16$ in our work). Dropout with rate $\delta=0.2$ is applied after each intermediate convolutional layer. The final class probabilities are obtained through:
$P(\mathbf{Y}|\mathbf{X})=\text{Softmax}(F_{N})$
Each convolutional layer captures local relationships between the predictions of both models, allowing the network to learn spatial patterns in prediction agreement or disagreement. This spatial-aware fusion can be particularly beneficial for complex scenes where the performance of individual models varies spatially.
IV-D Training
@@
$\mathcal{L}=-\sum_{n=1}^{N}\sum_{k=1}^{K}w_{k}y_{n,k}\log(\hat{y}_{n,k})$
where $K$ represents the number of classes, $w_{k}$ denotes the class weight assigned to class $k$, $y_{n,k}$ is the ground truth binary indicator (1 if sample $n$ belongs to class $k$, and 0 otherwise), and $\hat{y}_{n,k}$ represents the model’s predicted probability that sample $n$ belongs to class $k$. The class weights $w_{k}$ are computed as the inverse of the class frequencies in the training set, helping to address the inherent imbalance between cloud, shadow, and clear sky spatial soundings.
For optimization, we employ the Adam optimizer with a learning rate determined through cross-validation experiments detailed in Appendix -A. To enhance model generalization, we implement data augmentation techniques including random horizontal and vertical flips, and rotations at multiples of 90 degrees. These augmentations help the models learn invariance to common geometric transformations while preserving the physical meaning of the spectral signatures. Training proceeds over a maximum of 100 epochs with randomly shuffled batches of size 32. We employ an early stopping strategy to prevent overfitting, where training is terminated if the validation loss does not improve for 20 consecutive epochs (patience period), and we save the best-performing model checkpoint based on validation performance for final evaluation.
IV-E Post-processing and Evaluation
IV-E1 Post-processing for MethaneSAT
@@ IV-E2 Evaluation
  • Precision: Represents the model’s exactness in identifying clouds and shadows, calculated as the ratio of correctly predicted cloud/shadow spatial soundings to all spatial soundings predicted as cloud/shadow by the model. High precision indicates a low false positive rate, meaning when the model predicts a spatial sounding as cloud or shadow, it is likely to be correct. Precision = $\frac{\text{True Positives}}{\text{True Positives + False Positives}}$
  • Recall: Measures the completeness of the model in detecting actual clouds and shadows, calculated as the ratio of correctly predicted cloud/shadow spatial soundings to all ground truth cloud/shadow in the dataset. High recall indicates the model’s ability to identify most of the actual clouds and shadows. Recall = $\frac{\text{True Positives}}{\text{True Positives + False Negatives}}$
@@
We use macro-averaging, computing the precision, recall, and F1-score (i.e. we compute the metrics for each class independently and average them to report final results). The reported values include the test set mean performance and standard deviation in 3 cross-validated folds, ensuring a robust evaluation of model performance. See Appendix -A for details of the hyperparameter selection.
@@ V-A MethaneAIR
Detailed hyperparameter experiments can be found in Appendix -A, and comprehensive predictions for all models can be found in Appendix -C.
@@ VI Discussion
The baseline ILR model (accuracy: 73.81±4.05 for MethaneAIR) captures basic spectral features but generates highly fragmented predictions with significant noise. This purely spectral approach, while computationally efficient, fails to leverage the spatial relationships crucial for distinguishing between spectrally similar features like shadows and dark surfaces. The MLP model (accuracy: 82.49±2.24%) shows improved performance but still suffers from noise and fragmentation due to its pixel-wise classification approach, as clearly visible in Figures 8 and 11.
The U-Net architecture (accuracy: 88.26±0.45% for MethaneAIR, 78.73±3.23% for MethaneSAT) represents a substantial advancement, leveraging its encoder-decoder structure with skip connections to produce spatially coherent predictions. While U-Net exhibits a tendency toward over-smoothed boundaries around cloud formations and their shadows, this characteristic can be advantageous depending on the specific application requirements. On the other hand, our proposed SCAN model (accuracy: 86.51±2.90% for MethaneAIR, 80.33±3.43% for MethaneSAT) demonstrates superior ability in detecting accurate boundaries through its spectral attention mechanism, though it still exhibits some noise in spectrally complex regions.
Notably, SCAN outperforms U-Net on MethaneSAT despite lower MethaneAIR performance. We attribute this reversal to MethaneSAT’s coarser spatial resolution ($\sim 100\times 400$ m vs. $\sim 25$ m) creating more spectrally mixed boundary spatial soundings where wavelength-specific attention provides better discrimination than spatial convolution. The 6-7% cross-dataset performance gap underscores generalization challenges between sensors due to spatial resolution differences, scene complexity variations, and platform-specific characteristics. To validate SCAN’s design, we benchmarked against state-of-the-art attention mechanisms (Appendix -B). SCAN achieves 8.88-10.5 percentage point F1 improvements over SE-UNet, ViT-SegFormer, and SpectralFormer, demonstrating that spectral attention outperforms complex self-attention architectures when training data is limited (262 scenes).
@@
These findings have direct implications for the MethaneSAT mission, which requires accurate identification of atmospheric artifacts to ensure reliable methane retrievals. The superiority of the Combined CNN approach in handling complex scenes with varying surface reflectance is particularly relevant for global methane monitoring, where diverse terrain and atmospheric conditions can significantly impact detection accuracy. Our methods should generalize to other hyperspectral missions facing similar atmospheric preprocessing challenges, improving the reliability of methane plumes mapping and supporting global efforts to identify and quantify methane emissions across different platforms.
@@ VII Conclusions
From a computational perspective, our analysis demonstrates that the Combined CNN model offers an effective balance between performance and efficiency. While requiring moderately increased computational resources compared to individual models (4.1 ms inference time versus 1.7-2.1 ms per 1,000 km²), its significantly improved segmentation accuracy justifies this overhead for operational deployment in satellite missions where accurate atmospheric artifact detection is critical for reliable methane retrievals and subsequent emission quantification. Furthermore, the patch-based evaluation strategy demonstrated for MethaneSAT’s variable-sized inputs enhances the practical applicability of these methods in operational contexts.
Our findings have direct and urgent implications for global methane monitoring and climate change mitigation efforts. Accurate cloud and shadow masking is not merely a technical preprocessing: it is a critical prerequisite for reliable atmospheric CH4 retrievals that directly impact our ability to track progress toward international commitments in addressing climate change. Furthermore, our results offer immediately applicable insights and methodologies to other operational and planned hyperspectral missions facing similar atmospheric challenges in remote sensing-based greenhouse gas monitoring, amplifying the impact of this work across the growing constellation of methane monitoring capabilities.
@@ VIII Acknowledgments
Funding for MethaneSAT and MethaneAIR activities was provided in part by Anonymous, Arnold Ventures, The Audacious Project, Ballmer Group, Bezos Earth Fund, The Children’s Investment Fund Foundation, Heising-Simons Family Fund, King Philanthropies, Robertson Foundation, Skyline Foundation and Valhalla Foundation. For a more complete list of funders, please visit www.methanesat.org. We express our gratitude to the MethaneAIR field team for their contributions to MethaneAIR data collection. We thank the AstroAI and EarthAI institutes at the Center for Astrophyiscs $|$ Harvard & Smithsonian for useful discussions and guidance. CG and MPC were supported by AstroAI at the Center for Astrophysics $|$ Harvard and Smithsonian.